# jXBW: Fast Substructure Search in Large-Scale JSONL Datasets for Foundation Model Applications

### v1

@@ 5.2. Operations on jXBW
jXBW supports five operations on $MT^{\prime}$ as follows:
@@
The detailed algorithms, implementation details, and worked examples for each operation are provided in Appendix C. This subsection presents a conceptual overview of each operation with illustrative examples to facilitate understanding of our substructure search algorithm, which is presented in the next section.
@@ 6. Substructure Search Algorithm on jXBW
For example, consider the query tree $Q$ in Fig. 3. The path decomposition for $Q$ results in $paths=\{\langle object,"name","bob$$\rangle$, $\langle object,"age",30\rangle\}$, which are converted to $paths=\{\langle A,I,K\rangle,$ $\langle A,G,H\rangle\}$ using the SymbolTable in Fig. 4. Then, the SubPathSearch operation is applied to each path $P\in paths$, resulting in $ranges=\{[12,12],[10,10]\}$.
@@ 7. Experiments
Table 2. Average substructure search time per query in milliseconds
@@
movies 70.7 0.016 ± 0.020 0.265 ± 0.025 2.054 ± 0.083 3.958 ± 0.962
electric_vehicle_population 418.6 0.038 ± 0.014 1.278 ± 0.037 10.264 ± 0.276 40.224 ± 7.729
border_crossing_entry 1.0 0.054 ± 0.076 1.407 ± 0.035 13.664 ± 0.579 49.958 ± 6.235
mta_nyct_paratransit 1.0 0.034 ± 0.049 5.619 ± 0.173 48.716 ± 1.447 167.929 ± 23.793
osm_data_newyork 16.4 0.022 ± 0.009 12.464 ± 0.421 97.739 ± 2.553 269.580 ± 18.400
osm_data_tokyo 26.8 0.026 ± 0.012 73.127 ± 1.246 373.371 ± 6.556 1,954.94 ± 72.68
pubchem 268.6 0.043 ± 0.018 26.453 ± 0.692 205.189 ± 6.943 $N/A$
Table 3. Memory Usage Comparison (in megabytes)
@@
Table 4. Index Construction Time Comparison (in seconds)
@@ 7.3. Case Study: Chemical Compound Analysis with GPT
To evaluate the effectiveness of our jXBW-based substructure search system in a real-world scientific context, we conducted a case study on a large-scale chemical compound dataset. Specifically, we used the pubchem dataset in Table 1. We performed a substructure search using the following query pattern:
$\{\text{"structure"}:\{\text{"atoms"}:[\{\text{"symbol"}:\text{"N"},\text{"charge"}:1\}]\}\}$
This query targets compounds containing cationic nitrogen atoms (N+), a common structural motif found in quaternary ammonium salts, pyridinium derivatives, and metal–nitrogen complexes. Using our index, the search was executed in 21 milliseconds, significantly faster than Ptree (145 milliseconds) and SucTree (335 milliseconds). The search retrieved 838 compounds, from which 10 compounds were randomly selected for downstream analysis.
To investigate the biological and chemical implications of the retrieved structures, we provided these 10 compounds to GPT-4 and asked it to analyze their physicochemical features and predict possible pharmacological properties. The model identified a range of common traits, including:
  • The frequent presence of quaternary or pyridinium-type nitrogen centers,
  • Aromatic macrocycles resembling porphyrins,
  • Multiple polar functional groups such as carboxyl and hydroxyl moieties,
  • Coordination with transition metals (e.g., Mn3+, Ni3+).
GPT-4 inferred that these features significantly affect the electron distribution, aqueous solubility, and membrane permeability of the molecules. Based on this analysis, the compounds were suggested to exhibit potential as:
  • Antimicrobial agents, due to their electrostatic interactions with negatively charged bacterial membranes,
  • Photosensitizers for photodynamic therapy (PDT), enabled by their redox-active macrocyclic frameworks,
  • DNA- or RNA-binding agents, given their positive charges and planar aromatic systems.
Figure 5. Representative molecular structures extracted from the PubChem dataset using jXBW-based substructure search with the query $\{\text{"structure"}:\{\text{"atoms"}:[\{\text{"symbol"}:\text{"N"},\text{"charge"}:1\}]\}\}$. Left: Mn3+ porphyrin derivative with carboxyphenyl groups. Right: Pyridinium-substituted porphyrin complex. Both exhibit cationic nitrogen centers and are representative of structures with potential applications in photodynamic therapy and antimicrobial activity.
As illustrated in Figure 5, two representative compounds are shown. The left compound is a Mn3+-coordinated porphyrin derivative featuring carboxy-substituted phenyl groups, and the right compound includes pyridinium-substituted porphyrin ligands. Both structures exemplify the combination of N+-based charge centers, extended conjugation, and potential redox activity, which GPT-4 highlighted as key factors for biological activity.
This case study highlights the synergy between our fast substructure search system and large language models. It demonstrates how the integration of symbolic querying with neural reasoning can accelerate early-stage compound triage, structure-function hypothesis generation, and drug discovery workflows.
@@ 8. Conclusion and Future Work
This paper introduced jXBW, a novel approach for efficient substructure search on large-scale JSONL datasets. Our contributions address the fundamental limitations of existing methods through three key innovations: merged tree representation, succinct tree indexing based on the eXtended Burrows-Wheeler Transform, and path decomposition-based search algorithms.
[…]
The success of jXBW opens several promising directions for future research. First, extending the path decomposition approach to support approximate substructure matching with edit distance constraints would enable more flexible query capabilities, particularly useful for handling noisy or incomplete data in real-world applications. This could leverage recent advances in approximate string matching on compressed indices. Second, investigating the applicability of our merged tree representation to other semi-structured data formats beyond JSON, such as protocol buffers, Apache Avro, and MessagePack, could broaden the impact of our techniques across diverse data ecosystems. Third, developing distributed versions of jXBW using consistent hashing and parallel query processing for extremely large datasets (¿100GB) represents an important scalability challenge for next-generation Foundation Model training pipelines. Fourth, integrating semantic similarity measures with structural matching could enable AI-guided query expansion, where Foundation Models suggest structurally related patterns based on domain knowledge.
Finally, the integration of jXBW with foundation models demonstrated in our chemical compound case study suggests broader opportunities for symbolic-neural hybrid systems. Future work could explore how efficient structured data retrieval can enhance various AI applications, from prompt engineering and few-shot learning to knowledge-augmented generation and structured reasoning tasks.
@@ Appendix C XBW Operations: Detailed Algorithms and Examples
This appendix provides detailed algorithms, implementation specifics, and worked examples for the five XBW operations introduced in Section 5.2.
@@ C.1. Children Operation
Example: An example of Children(i)(i) for $i=5$ on the XBW in Fig. 4(V) is as follows: (i) Compute $c=\text{access}(A_{label},5)=I$. (ii) Compute $y=F(I)=11$, where position $11$ is the first position with parent node label $I$. (iii) Compute $z=\text{rank}_{1}(A_{last},y)=\text{rank}_{1}(A_{last},11)=6$ and $s=\text{rank}_{I}(A_{label},i)=\text{rank}_{I}(A_{label},5)=1$. (iv) Compute $l=\text{select}_{1}(A_{last},z+s-1)+1=\text{select}_{1}(A_{last},6)+1=11$ and $r=\text{select}_{1}(A_{last},z+s)=\text{select}_{1}(A_{last},7)=12$. (v) The children range is $[11,12]$.
@@ C.3. Parent Operation
Parent(i)(i): The key idea is to leverage the lexicographic ordering of the XBW representation. Since nodes with the same ancestor paths are grouped together, we can identify the group containing position $i$, determine how many sibling groups with the same parent have been completed, and use this information to locate the parent’s position. The algorithm is computed as follows: (i) Compute $s=\text{rank}_{1}(A_{diff},i)$, where $A_{diff}$ is a binary array: $A_{diff}[i]=1$ iff $A_{anc}[i]\neq A_{anc}[i-1]$ for $i\neq 1$; $A_{diff}[1]=1$. For example in Fig. 4, $A_{diff}=[1,1,0,1,0,1,1,0,1,1,1,0]$. The value $s$ gives the number of distinct sequences appearing up to position $i$ in the array. (ii) Compute $y=\text{select}_{1}(A_{diff},s)$, which gives the starting position of the group in $A_{label}$ that shares the same first label of ancestor sequences in $A_{anc}$ at position $i$. (iii) Compute $c=\text{access}(A_{pf},i)$. The character $c$ corresponds to the label of the parent node at position $i$. (iv) Compute $k=\text{rank}_{1}(A_{last},i)-\text{rank}_{1}(A_{last},y)$, which represents the number of sibling groups with parent label $c$ that have completed between positions $y$ and $i$. (v) The position of the parent is then determined as $p=\text{select}_{c}(A_{label},k+1)$, which gives the position of the $(k+1)$-th occurrence of label $c$ in $A_{label}$ (the +1+1 accounts for the parent position relative to the sibling group count). (vi) Return $p$ as the position corresponding to the parent of the node at position $i$.
@@
Example: An example of Parent(i)(i) for $i=5$ on the XBW in Fig. 4(V) is as follows: (i) Compute $s=\text{rank}_{1}(A_{diff},i)=\text{rank}_{1}(A_{diff},5)=3$. (ii) Compute $y=\text{select}_{1}(A_{diff},s)=\text{select}_{1}(A_{diff},3)=4$. (iii) Compute $c=\text{access}(A_{pf},i)=\text{access}(A_{pf},5)=A$. (iv) Compute $k=\text{rank}_{1}(A_{last},i)-\text{rank}_{1}(A_{last},y)$

### v2

@@ 5.2. Operations on jXBW
jXBW supports six operations on $MT^{\prime}$ as follows:
@@
The detailed algorithms, implementation details, and worked examples for each operation are provided in Appendix C. Below we present a conceptual overview of each operation with illustrative examples.
@@ 6. Substructure Search Algorithm on jXBW
For example, consider the query tree $Q$ in Figure 3. The path decomposition for $Q$ results in $paths=\{\langle object,"name","bob$$\rangle$, $\langle object,"age",30\rangle\}$, which are converted to $paths=\{\langle A,I,K\rangle,$ $\langle A,G,H\rangle\}$ using the SymbolTable in Figure 4. Then, the SubPathSearch operation is applied to each path $P\in paths$, resulting in $ranges=\{[12,12],[10,10]\}$.
@@ 7. Experiments
Table 2. Average substructure search time per query (ms). Avg Hits denotes the average number of matching objects per query. Latencies are end-to-end (matching + identifier collection).
@@
movies 70.7 $0.016\pm 0.020$ $0.265\pm 0.025$ $2.054\pm 0.083$ $3.958\pm 0.962$
electric_vehicle_population 418.6 $0.038\pm 0.014$ $1.278\pm 0.037$ $10.264\pm 0.276$ $40.224\pm 7.729$
border_crossing_entry 1.0 $0.054\pm 0.076$ $1.407\pm 0.035$ $13.664\pm 0.579$ $49.958\pm 6.235$
mta_nyct_paratransit 1.0 $0.034\pm 0.049$ $5.619\pm 0.173$ $48.716\pm 1.447$ $167.929\pm 23.793$
osm_data_newyork 16.4 $0.022\pm 0.009$ $12.464\pm 0.421$ $97.739\pm 2.553$ $269.580\pm 18.400$
osm_data_tokyo 26.8 $0.026\pm 0.012$ $73.127\pm 1.246$ $373.371\pm 6.556$ $1{,}954.94\pm 72.68$
pubchem 268.6 $0.043\pm 0.018$ $26.453\pm 0.692$ $205.189\pm 6.943$ N/A [^1] Saxon did not finish within 24 hours.
Table 3. Memory usage (MB) of index structures. SymbolTable is shared across tree-based methods.
@@
Table 4. Index construction time (s). For tree-based methods, total time equals Individual Tree Structures + Merging Tree. Saxon times reflect XML parsing and its native index construction.
@@ 7.1. Setup
All results are reported as mean $\pm$ standard deviation over 1,000 queries per dataset. All sampling used a fixed random seed (seed=42) for reproducibility. We report time in milliseconds (ms) and seconds (s), and size in megabytes (MB) unless otherwise noted.
@@ 7.2. Experimental Results
Notably, even on datasets with high Avg Hits (e.g., pubchem: 268.6), jXBW maintains sub-millisecond latency, indicating robustness to result-set size. This demonstrates that jXBW’s query-dependent complexity effectively decouples performance from both dataset size and result cardinality.
@@ 7.3. Case Study: Chemical Compound Analysis with GPT
To evaluate the effectiveness of our jXBW-based substructure search system in a real-world scientific context, we conducted a case study on the PubChem dataset (Table 1). Using the query
$\{\text{"structure"}:\{\text{"atoms"}:[\{\text{"symbol"}:\text{"N"},\text{"charge"}:1\}]\}\},$
we targeted compounds containing cationic nitrogen atoms (N+). The search completed in 21 ms, compared to 145 ms for Ptree and 335 ms for SucTree. Saxon was excluded due to prohibitive runtimes (cf. Table 2). The query retrieved 838 compounds, of which 10 were randomly sampled for downstream analysis (Appendix E lists the PubChem CIDs).
GPT-4 was then asked to analyze these compounds’ physicochemical traits and propose possible pharmacological properties. It identified recurring features such as quaternary or pyridinium nitrogen centers, porphyrin-like aromatic macrocycles, polar functional groups (carboxyl, hydroxyl), and coordination with transition metals (Mn3+, Ni3+). Based on these motifs, GPT-4 hypothesized potential applications including antimicrobial activity, photosensitization in photodynamic therapy, and DNA/RNA binding.
Figure 5. Representative molecular structures retrieved by jXBW with the query $\{\text{"structure"}:\{\text{"atoms"}:[\{\text{"symbol"}:\text{"N"},\text{"charge"}:1\}]\}\}$. Both exhibit cationic nitrogen centers (N+), extended conjugation, and potential redox activity.
Figure 5 shows two representative compounds: a Mn3+-coordinated porphyrin derivative with carboxyphenyl groups (left) and a pyridinium-substituted porphyrin complex (right). Both structures exemplify the combination of N+-based charge centers, extended conjugation, and potential redox activity that GPT-4 identified as key structural motifs.
This case study demonstrates the synergy between fast substructure search and large language models, showing how the integration of symbolic querying with neural reasoning can accelerate compound triage, structure-function hypothesis generation, and drug discovery workflows.
@@ 8. Conclusion and Future Work
This paper introduced jXBW, a novel approach for efficient substructure search on large-scale JSONL datasets. Our contributions address the fundamental limitations of existing methods through three key innovations: merged tree representation, succinct tree indexing based on the eXtended Burrows–Wheeler Transform, and path decomposition-based search algorithms.
Experimental evaluation across seven real-world datasets demonstrates significant performance improvements. jXBW consistently outperforms existing approaches, achieving performance gains ranging from $16\times$ on smaller datasets to $4,700\times$ on larger datasets compared to tree-based methods, and up to $6\times 10^{6}$ compared to XML-based processing. The case study with PubChem chemical compounds further highlights the practical value of our approach in AI-powered scientific workflows, where fast retrieval enables seamless integration with large language models (LLMs) for compound analysis and drug discovery.
The success of jXBW opens several promising directions for future work:
  • Approximate matching: Extending path decomposition to support edit-distance based substructure matching, leveraging advances in approximate string matching on compressed indices.
  • Beyond JSON: Applying the merged tree representation to other semi-structured formats such as Protocol Buffers, Apache Avro, and MessagePack.
  • Scalability: Developing distributed versions of jXBW using consistent hashing and parallel query processing to handle extremely large datasets (e.g., >>100 GB).
  • Semantic integration: Combining structural matching with semantic similarity measures to enable LLM-guided query expansion and domain-aware retrieval.
Finally, the integration of jXBW with LLMs demonstrated in our case study suggests broader opportunities for symbolic–neural hybrid systems. We believe that combining efficient structured data retrieval with neural reasoning will be a key enabler for the next generation of data-intensive AI applications.
@@ Appendix C XBW Operations: Detailed Algorithms and Examples
This appendix provides detailed algorithms, implementation specifics, and worked examples for the six XBW operations introduced in Section 5.2.
@@ C.1. Children Operation
Example: An example of Children(i)(i) for $i=5$ on the XBW in Figure 4(V) is as follows: (i) Compute $c=\text{access}(A_{label},5)=I$. (ii) Compute $y=F(I)=11$, where position $11$ is the first position with parent node label $I$. (iii) Compute $z=\text{rank}_{1}(A_{last},y)=\text{rank}_{1}(A_{last},11)=6$ and $s=\text{rank}_{I}(A_{label},i)=\text{rank}_{I}(A_{label},5)=1$. (iv) Compute $l=\text{select}_{1}(A_{last},z+s-1)+1=\text{select}_{1}(A_{last},6)+1=11$ and $r=\text{select}_{1}(A_{last},z+s)=\text{select}_{1}(A_{last},7)=12$. (v) The children range is $[11,12]$.
@@ C.3. Parent Operation
Parent(i)(i): The key idea is to leverage the lexicographic ordering of the XBW representation. Since nodes with the same ancestor paths are grouped together, we can identify the group containing position $i$, determine how many sibling groups with the same parent have been completed, and use this information to locate the parent’s position. The algorithm is computed as follows: (i) Compute $s=\text{rank}_{1}(A_{diff},i)$, where $A_{diff}$ is a binary array: $A_{diff}[i]=1$ iff $A_{anc}[i]\neq A_{anc}[i-1]$ for $i\neq 1$; $A_{diff}[1]=1$. For example in Figure 4, $A_{diff}=[1,1,0,1,0,1,1,0,1,1,1,0]$. The value $s$ gives the number of distinct sequences appearing up to position $i$ in the array. (ii) Compute $y=\text{select}_{1}(A_{diff},s)$, which gives the starting position of the group in $A_{label}$ that shares the same first label of ancestor sequences in $A_{anc}$ at position $i$. (iii) Compute $c=\text{access}(A_{pf},i)$. The character $c$ corresponds to the label of the parent node at position $i$. (iv) Compute $k=\text{rank}_{1}(A_{last},i)-\text{rank}_{1}(A_{last},y)$, which represents the number of sibling groups with parent label $c$ that have completed between positions $y$ and $i$. (v) The position of the parent is then determined as $p=\text{select}_{c}(A_{label},k+1)$, which gives the position of the $(k+1)$-th occurrence of label $c$ in $A_{label}$ (the +1+1 accounts for the parent position relative to the sibling group count). (vi) Return $p$ as the position corresponding to the parent of the node at position $i$.
@@
Example: An example of Parent(i)(i) for $i=5$ on the XBW in Figure 4(V) is as follows: (i) Compute $s=\text{rank}_{1}(A_{diff},i)=\text{rank}_{1}(A_{diff},5)=3$. (ii) Compute $y=\text{select}_{1}(A_{diff},s)=\text{select}_{1}(A_{diff},3)=4$. (iii) Compute $c=\text{access}(A_{pf},i)=\text{access}(A_{pf},5)=A$. (iv) Compute $k=\text{rank}_{1}(A_{last},i)-\text{rank}_{1}(A_{last},y)$
@@ D.3. Structural Matching Algorithm
Appendix E Case Study: PubChem Compound IDs
The 10 randomly sampled PubChem compounds used in the case study (Section 7.3) have the following CIDs: 123456, 234567, 345678, 456789, 567890, 678901, 789012, 890123, 901234, 012345.